what is 13= w/-3 -4? and what is -17=-32-3/5f?

Answer by tutorcecilia(2152)

13= w/-3 -4 [Solve for w by isolating the w]
13+w=2/-3-4+4
17=w/-3
(17)(-3)=(w/-3)(-3/1)
-51=w
.
Check by plugging (w=-51) back into the original equatin]
13= w/-3 -4
13= (-51)/-3 -4
13=13 [checks out]
.
-17=-32-3/5f [Solve for f]
-17+32=-32+32-3/5f
-17+32=-32+32-3/5f
15=-3/5f
(15)(5f/1)=(-3/5f)(5f/1)
75f=-3
(1/75)(75f)=(1/75)(-3)
f=-3/75
f=-1/25
.
Check by plugging (f=-1/25) back into the original equation